# Viscoplasticity can stabilise liquid collar motion on vertical cylinders

### v2

@@ Appendix A Small-$B$ asymptotics in a low-yield-stress regime
A good approximation for $I_{0}$ is given by $VI_{0}\approx 8.9\Gamma$, at least when $200<\Gamma<$ 20 (figure 8a). For the values of $B$ used in figure 5, i.e., for $B\geq 0.025$, $\Gamma<20$ holds for $0\leq\mathcal{J}\leq 0.4$. Substituting the linear approximation, $I_{0}\approx 8.9\Gamma$, into (A.14), we find
@@
Figure 8: Solution to (A.12) with (A.13) and (A.2) (black). Red dashed lines show the linear approximation, $VI_{0}\approx 8.9\Gamma$. (a) $VI_{0}$ is approximated well by $8.9\Gamma$ for $200<\Gamma<$ 20. (b) At large values of $\Gamma$, the slope of $VI_{0}$ approaches $2\pi$.
Figure 5 suggests that the collar speed, $\mathcal{U}$, determined via (A.15), provides improved agreement with numerical simulations at small values of $\mathcal{J}=B^{1/2}\hat{J}$, compared to the speed prediction from (76). Since the values of $B$ and $J$ used in simulations in figure 5 are not extremely small, it is not necessarily clear whether the assumption that $J=O(B)$ made in §IV, or the assumption that $J=O(B^{3/2})$ made in this appendix, or another scaling for $J$, is most appropriate for each simulation. Figure 9 shows the locations of each of those simulations from figure 5 in $J$-$B$ space, along with the lines $J=B$ and $J=B^{3/2}$. We can expect those simulations lying close to $J=B$ may obey the $J=O(B)$ asymptotic regime, and those lying close to the line $J=B^{3/2}$ may obey the $J=O(B^{3/2})$ asymptotic regime more closely. However, figure 9 indicates that at those values of $B$ used in simulations, there is generally not a great deal of separation between $B$ and $B^{3/2}$, so for many of the simulations it is not immediately clear which regime may provide the better predictions. We can expect that, if much smaller values of $B$ were able to be accessed, simulations with parameter values near the $J=B$ line would obey the $J=O(B)$ theory closely, simulations with parameter values near the line $J=B^{3/2}$ would obey the $J=O(B^{3/2})$ theory closely, and those with parameter values lying far below the line $J=B^{3/2}$ would exhibit essentially Newtonian behaviour.
@@ Appendix B Inner regions within the front Bretherton region
In the small-$B$ asymptotic model presented in §IV, the leading-order layer height in the front Bretherton region has an undulating shape, with $\mathcal{H}(\xi)=H_{\infty}^{-}$ at a series of points, which we shall call $\xi=\{\xi_{0},\xi_{1},\xi_{2},\dots\}$, with $\xi_{0}<\xi_{1}<\dots$. Sufficiently close to each of these points, the expansions (($)61𝑎𝑏$) are no longer valid, as the final term in the yield criterion (32) enters the leading-order balance. To resolve this, we introduce short inner regions, of length $O(B)$, around each point, $\xi=\{\xi_{0},\xi_{1},\xi_{2},\dots\}$, in which $h\approx BH_{\infty}^{-}$ and $Y$ is not necessarily close to $h$. We expect that, within these inner regions, there will be only small changes in $h$ compared to the ‘outer’ Bretherton region solution, i.e. the solution of (62), but that there may be significant variation in $Y$ within these regions, reflecting the fact that the leading-order shear stress changes sign across each inner region.
@@
where $\mathcal{H}_{-}(\xi)$ is the solution to (62). In the neighbourhood of $\xi=\xi_{n}$, i.e. where $Z=O(1)$ and $|\mathcal{H}_{-}-H_{\infty}^{-}|=O(|\xi-\xi_{n}|)$, (62) implies that ′′′$\mathcal{H}_{-}^{\prime\prime\prime}=O(|\xi-\xi_{n}|)$, so
@@
Inserting (B.1) and (B.2) into (30) gives
$\frac{\mathcal{U}\alpha_{n}Z}{H_{\infty}^{-}-\tilde{y}/3}=\frac{1}{2}\tilde{y}^{2}\left(\frac{3\mathcal{U}\alpha_{n}Z}{(H_{\infty}^{-})^{3}}+\tilde{H}^{\prime\prime\prime}\right),$ (B.3)
@@
$\mathcal{J}=(H_{\infty}^{-}-\tilde{y})\left|\frac{3\mathcal{U}\alpha_{n}Z}{(H_{\infty}^{-})^{3}}+\tilde{H}^{\prime\prime\prime}\right|.$ (B.4)
@@
$\mathcal{J}\tilde{y}^{2}\left(H_{\infty}^{-}-\frac{\tilde{y}}{3}\right)=|2\mathcal{U}\alpha_{n}Z|(H_{\infty}^{-}-\tilde{y}),$ (B.5)
@@
We expect that there is not necessarily an infinite sequence of inner regions for which this solution holds. Rather, we expect there to be a finite distance at which the Bretherton region solution, outlined in §IV.4, and hence the inner-region solutions outlined in this section, break down. We expect this to occur when $|\xi|=O(\log({1/B}))$ with $\xi<0$, at which point $|H_{\infty}^{-}-\mathcal{H}_{-}|=O(B^{1/2})$ (this can be seen from (62) and the exponential decay of $\mathcal{H}_{-}$ towards $H_{\infty}^{-}$ in the far-field) and then the assumption in (($)61𝑎𝑏$) that $Y\approx h$ no longer holds. Therefore, we expect there to be a finite sequence of the inner regions we have outlined here, around points $\xi=\{\xi_{0},\xi_{1},\dots,\xi_{N}\}$, for some $N$, before the Bretherton solution breaks down and the fluid transitions towards becoming unyielded. In appendix C, we outline how the layer transitions from being in an almost fully-yielded state in the Bretherton regions (both at the front and rear of the collar) to becoming unyielded at a finite distance from the collar.
@@ Appendix C Rigidification regions
The solutions, $H_{\pm}$, to (63) decay exponentially towards $1$ as $X\rightarrow\pm\infty$ in the far fields away from the collar. We expect that the Bretherton region solutions, outlined in §IV.4, will break down when when |±$|\mathcal{H}_{\pm}-H_{\infty}^{\pm}|=O(B^{1/2})$, as we would no longer expect that $Y\approx B\mathcal{H}_{\pm}$ to leading order, as is assumed in (($)61𝑎𝑏$). We have identified this breakdown of the Bretherton region solution in multiple short inner regions in §B, but we expect that it also occurs in other regions close to the points at which the film thickness becomes uniform and $Y\rightarrow 0^{+}$, at the front and rear ends of the collar. We call these regions ‘rigidification regions’, since within them, $Y$ transitions from being close to $h$ to being close to zero. We outline the structures of these regions, at the front and rear of the collar, below.
At the rear, we suppose that there is a point, $\zeta=\zeta_{0}$, such that $Y\rightarrow 0^{+}$ as $\zeta\rightarrow\zeta_{0}^{-}$ (figure 3c). We then define a local coordinate, $\tilde{z}$, such that $\zeta=\zeta_{0}-B^{1/2}\tilde{z}$. If we make the expansions (($)61𝑎𝑏$), we recover the classical equation (62) from the Bretherton region. When $h=B(H_{\infty}^{+}+\tilde{H})$, where $\tilde{H}\ll H_{\infty}^{+}$, (62) then suggests exponential dependence of $\tilde{H}$ on $\tilde{z}$. However, from (32), it can be seen that the assumption that $Y\approx h$ is violated when $\tilde{H}=O(B^{1/2})$, which, given the exponential dependence of $\tilde{H}$ on $\tilde{z}$, occurs when $\tilde{z}=O(\log({1/B}))$. Therefore, strictly, we expect the solution outlined in §IV.4 to hold when $|\tilde{z}|\gg\log({1/B})$, and we require a separate inner solution when $|\tilde{z}|\ll\log({1/B})$. This motivates the $O(B^{1/2}\log(1/B))$ length scale identified in the sketch in figure 3(c) for the distance between $\zeta=\zeta_{0}$ and the Bretherton region. We call the inner region where $\tilde{z}\ll\log(1/B)$ a ‘rigidification region’.
@@
$h=BH_{\infty}^{+}+B^{3/2}\tilde{H}_{+}(\tilde{z})+\dots,\quad Y=B\tilde{Y}_{+}(\tilde{z}).$ (C.1)
@@
$-\frac{1}{2}\tilde{Y}_{+}^{2}\tilde{H}_{+,\tilde{z}\tilde{z}\tilde{z}}=\frac{\mathcal{U}\tilde{H}_{+}}{H_{\infty}^{+}-\frac{\tilde{Y}_{+}}{3}},$ (C.2)
@@
$h=BH_{\infty}^{-}+B^{3/2}\tilde{H}_{-}(\tilde{z})+\dots,\quad Y=B\tilde{Y}_{-}(\tilde{z}),$ (C.5)
@@
Again, rather than solving the equations (C.6) and (C.7) in full, we focus on their inner and outer limits. In the outer limit of the rigidification region, where $|\tilde{z}|\rightarrow\infty$, we again recover the exponential dependence of $\tilde{H}$ on $\tilde{z}$, so that this solution can, in theory, be matched onto the undulating solution for $h$ in the Bretherton region. However, at the front, we do not expect the inner limit of the rigidification region to have the structure (C.4). Instead, we expect a more intricate structure in the rigidification region, with $\tilde{H}$ potentially undulating and repeatedly changing signs. In the inner limit, where $\tilde{z}\rightarrow 0^{+}$ and $\tilde{Y}\rightarrow 0^{+}$, (C.7) reduces to
@@
where we have also used (C.6), and the fact that $\mathcal{U}\leq 0$, to reach (C.8). The structure of the solution to (C.8) was detailed by Jalaal et al. [39], since the same equation arises at the leading edge of a droplet spreading into a uniform precursor film. The solution for $\tilde{H}_{-}$ is composed of an infinite sequence of cubic polynomials, with the $n^{\mathrm{th}}$ cubic being defined in an interval, $I_{n}=[\tilde{z}_{n},\tilde{z}_{n+1}]$, and continuity of $\tilde{H}_{-}$ and its first and second derivatives being enforced at the boundaries between each of the intervals $\{I_{n}\}$. The third derivative is not, in general, continuous at $\tilde{z}=\tilde{z}_{n}$. This results in an undulating structure for $\tilde{H}_{-}$, with both the amplitude and wavelength of the undulations decreasing as $n\rightarrow\infty$, with the train of undulations ending at a finite distance [39], which we can identify as the point $\tilde{z}=0$. Again, this structure is inconsistent (over very small scales) with the long-wave assumption underpinning the original model, but we do not pursue this detail here.
@@ Appendix D Higher-order corrections to the small-$B$ asymptotic solution
$h$ $=$ $B^{1/2}\frac{V}{4\pi}\hat{X}^{2}+B\left(\hat{A}_{\pm}-\frac{V\hat{X}^{4}}{48\pi}\right)+B^{5/4}\hat{h}_{\pm}+\dots,$ (D.9)
$Y$ $=$ $B^{1/2}\hat{Y}_{\pm}+\dots,$ (D.10)
where $\hat{h}_{\pm}(\hat{X})$ and $\hat{Y}_{\pm}(\hat{X})$ are determined via (54) and (55). In the limits of the intermediate regions approaching the main collar body, (54) and (55) give
@@
$-\frac{V\hat{X}^{2}}{4\pi}\hat{h}_{\pm}^{\prime\prime\prime}\sim\mathcal{J}+\frac{1}{|\hat{X}|}\sqrt{\frac{8\pi|\mathcal{U}|\mathcal{J}}{V}}+\dots\quad\mbox{as}\quad\hat{X}\rightarrow\mp\infty,$ (D.14)
@@
We can approximate $D_{1}$ by using the expression for collar speed determined from leading-order theory (76), then determine the adjusted expression for $\mathcal{U}$ via (D.18); this results in generally improved agreement between asymptotic predictions for collar speeds and collar speeds in numerical simulations (figure 5).

### v3

@@ Appendix A Small-$B$ asymptotics in a low-yield-stress regime
A good approximation for $I_{0}$ is given by $VI_{0}\approx 8.9\Gamma$, at least when $0<\Gamma<20$ (figure 8a). For $B\geq 0.025$, $\Gamma<20$ holds for $0\leq\mathcal{J}\leq 0.4$. Substituting the linear approximation, $I_{0}\approx 8.9\Gamma$, into (A.14), we find
@@
Figure 8: Solution to (A.12) with (A.13) and (A.2) (black). Red dashed lines show the linear approximation, $VI_{0}\approx 8.9\Gamma$. (a) $VI_{0}$ is approximated well by $8.9\Gamma$ for $0<\Gamma<20$. (b) At large values of $\Gamma$, the slope of $VI_{0}$ approaches $2\pi$.
Figure 5(a) suggests that the collar speed, $\mathcal{U}$, determined via (A.15), provides improved agreement with numerical simulations at small values of $\mathcal{J}=B^{1/2}\hat{J}$, compared to the speed prediction from (76). Since the values of $B$ and $J$ used in simulations in figure 5 are not extremely small, it is not necessarily clear whether the assumption that $J=O(B)$ made in §IV, or the assumption that $J=O(B^{3/2})$ made in this appendix, or another scaling for $J$, is most appropriate for each simulation. Figure 9 shows the locations of each of those simulations from figure 5 in $J$-$B$ space, along with the lines $J=B$ and $J=B^{3/2}$. We can expect those simulations lying close to $J=B$ may obey the $J=O(B)$ asymptotic regime, and those lying close to the line $J=B^{3/2}$ may obey the $J=O(B^{3/2})$ asymptotic regime more closely. However, figure 9 indicates that at those values of $B$ used in simulations, there is generally not a great deal of separation between $B$ and $B^{3/2}$, so for many of the simulations it is not immediately clear which regime may provide the better predictions. We can expect that, if much smaller values of $B$ were able to be accessed, simulations with parameter values near the $J=B$ line would obey the $J=O(B)$ theory closely, simulations with parameter values near the line $J=B^{3/2}$ would obey the $J=O(B^{3/2})$ theory closely, and those with parameter values lying far below the line $J=B^{3/2}$ would exhibit essentially Newtonian behaviour.
@@ Appendix B Inner regions within the front Bretherton region
In the small-$B$ asymptotic model presented in §IV, the leading-order layer height in the front Bretherton region has an undulating shape, with $\mathcal{H}(\xi)=H_{\infty}^{-}$ at a series of points, which we shall call $\xi=\{\xi_{0},\xi_{1},\xi_{2},\dots\}$, with $\xi_{0}<\xi_{1}<\dots$. Sufficiently close to each of these points, the expansions ((61a,b)) are no longer valid, as the final term in the yield criterion (32) enters the leading-order balance. To resolve this, we introduce short inner regions, of length $O(B)$, around each point, $\xi=\{\xi_{0},\xi_{1},\xi_{2},\dots\}$, in which $h\approx BH_{\infty}^{-}$ and $Y$ is not necessarily close to $h$. We expect that, within these inner regions, there will be only small changes in $h$ compared to the ‘outer’ Bretherton region solution, i.e. the solution of (62), but that there may be significant variation in $Y$ within these regions, reflecting the fact that the leading-order shear stress changes sign across each inner region.
@@
where $\mathcal{H}_{-}(\xi)$ is the solution to (62). In the neighbourhood of $\xi=\xi_{n}$, i.e. where $Z=O(1)$ and $|\mathcal{H}_{-}-H_{\infty}^{-}|=O(|\xi-\xi_{n}|)$, (62) implies that $\mathcal{H}_{-}^{\prime\prime\prime}=O(|\xi-\xi_{n}|)$, so
$\mathcal{H}_{-}=H_{\infty}^{-}+\alpha_{n}(\xi-\xi_{n})+\frac{\beta_{n}}{2}(\xi-\xi_{n})^{2}+\frac{\mathcal{U}\alpha_{n}}{8(H_{\infty}^{-})^{3}}(\xi-\xi_{n})^{4}+\dots{\\ }=H_{\infty}^{-}+\alpha_{n}B^{1/2}Z+\frac{\beta_{n}}{2}BZ^{2}+\frac{\mathcal{U}\alpha_{n}}{8(H_{\infty}^{-})^{3}}B^{2}Z^{4}+\dots.$ (B.2)
Inserting (B.1) and (B.2) into (30) gives
@@
[…]
@@
[…]
@@
We expect that there is not necessarily an infinite sequence of inner regions for which this solution holds. Rather, we expect there to be a finite distance at which the Bretherton region solution, outlined in §IV.4, and hence the inner-region solutions outlined in this section, break down. We expect this to occur when $|\xi|=O(\log({1/B}))$ with $\xi<0$, at which point $|H_{\infty}^{-}-\mathcal{H}_{-}|=O(B^{1/2})$ (this can be seen from (62) and the exponential decay of $\mathcal{H}_{-}$ towards $H_{\infty}^{-}$ in the far-field) and then the assumption in ((61a,b)) that $Y\approx h$ no longer holds. Therefore, we expect there to be a finite sequence of the inner regions we have outlined here, around points $\xi=\{\xi_{0},\xi_{1},\dots,\xi_{N}\}$, for some $N$, before the Bretherton solution breaks down and the fluid transitions towards becoming unyielded. In appendix C, we outline how the layer transitions from being in an almost fully-yielded state in the Bretherton regions (both at the front and rear of the collar) to becoming unyielded at a finite distance from the collar.
@@ Appendix C Rigidification regions
The solutions, $H_{\pm}$, to (63) decay exponentially towards $1$ as $X\rightarrow\pm\infty$ in the far fields away from the collar. We expect that the Bretherton region solutions, outlined in §IV.4, will break down when when $|\mathcal{H}_{\pm}-H_{\infty}^{\pm}|=O(B^{1/2})$, as we would no longer expect that $Y\approx B\mathcal{H}_{\pm}$ to leading order, as is assumed in ((61a,b)). We have identified this breakdown of the Bretherton region solution in multiple short inner regions in §B, but we expect that it also occurs in other regions close to the points at which the film thickness becomes uniform and $Y\rightarrow 0^{+}$, at the front and rear ends of the collar. We call these regions ‘rigidification regions’, since within them, $Y$ transitions from being close to $h$ to being close to zero. We outline the structures of these regions, at the front and rear of the collar, below.
At the rear, we suppose that there is a point, $\zeta=\zeta_{0}$, such that $Y\rightarrow 0^{+}$ as $\zeta\rightarrow\zeta_{0}^{-}$ (figure 3c). We then define a local coordinate, $\tilde{z}$, such that $\zeta=\zeta_{0}-B^{1/2}\tilde{z}$. If we make the expansions ((61a,b)), we recover the classical equation (62) from the Bretherton region. When h=B(H++H~)h=B($_{\infty}^{+}+\tilde{H}$), where $\tilde{H}\ll H_{\infty}^{+}$, (62) then suggests exponential dependence of $\tilde{H}$ on $\tilde{z}$. However, from (32), it can be seen that the assumption that $Y\approx h$ is violated when $\tilde{H}=O(B^{1/2})$, which, given the exponential dependence of $\tilde{H}$ on $\tilde{z}$, occurs when $\tilde{z}=O(\log({1/B}))$. Therefore, strictly, we expect the solution outlined in §IV.4 to hold when $|\tilde{z}|\gg\log({1/B})$, and we require a separate inner solution when $|\tilde{z}|\ll\log({1/B})$. This motivates the $O(B^{1/2}\log(1/B))$ length scale identified in the sketch in figure 3(c) for the distance between $\zeta=\zeta_{0}$ and the Bretherton region. We call the inner region where $\tilde{z}\ll\log(1/B)$ a ‘rigidification region’.
@@
[…]
@@
$\frac{1}{2}\tilde{Y}_{+}^{2}\tilde{H}_{+,\tilde{z}\tilde{z}\tilde{z}}=\frac{\mathcal{U}\tilde{H}_{+}}{H_{\infty}^{+}-\frac{\tilde{Y}_{+}}{3}},$ (C.2)
@@
[…]
@@
[…]
@@
where we have also used (C.6), and the fact that $\mathcal{U}\leq 0$, to reach (C.8). The structure of the solution to (C.8) was detailed by Jalaal et al. [38], since the same equation arises at the leading edge of a droplet spreading into a uniform precursor film. The same equation also appears when considering long bubbles translating in tubes, providing the solution very close to where the fluid rigidifies in the uniform film deposited on the tube wall [39]. The difference in the present analysis is that, since we focus on the small-$B$ limit with $\mathcal{J}=O(1)$, the rigidification region is always adjacent to an almost fully-yielded ‘Bretherton’ region.
As explained by Jalaal et al. [38], the solution for $\tilde{H}_{-}$ in (C.8) is composed of an infinite sequence of cubic polynomials, with the $n^{\mathrm{th}}$ cubic being defined in an interval, $I_{n}=[\tilde{z}_{n},\tilde{z}_{n+1}]$, and continuity of $\tilde{H}_{-}$ and its first and second derivatives being enforced at the boundaries between each of the intervals $\{I_{n}\}$. The third derivative is not, in general, continuous at $\tilde{z}=\tilde{z}_{n}$. This results in an undulating structure for $\tilde{H}_{-}$, with both the amplitude and wavelength of the undulations decreasing as $n\rightarrow\infty$, with the train of undulations ending at a finite distance [38], which we can identify as the point $\tilde{z}=0$. Again, this structure is inconsistent (over very small scales) with the long-wave assumption underpinning the original model, but we do not pursue this detail here.
@@ Appendix D Higher-order corrections to the small-$B$ asymptotic solution
$h$ $=$ $B^{1/2}\frac{V}{4\pi}\hat{X}^{2}+B\left(\hat{A}_{\pm}-\frac{V\hat{X}^{4}}{48\pi}\right)+B^{5/4}\hat{h}_{\pm}$$\dots$, (D.9)
$Y$ $=$ $B^{1/2}\hat{Y}_{\pm}$$\dots$, (D.10)
[…]
@@
$\frac{V\hat{X}^{2}}{4\pi}\hat{h}_{\pm}^{\prime\prime\prime}\sim\mathcal{J}+\frac{1}{|\hat{X}|}\sqrt{\frac{8\pi|\mathcal{U}|\mathcal{J}}{V}}+\dots\quad\mbox{as}\quad\hat{X}\rightarrow\mp\infty$, (D.14)
@@
We can approximate $D_{1}$ by using the expression for collar speed determined from leading-order theory (76), then determine the adjusted expression for $\mathcal{U}$ via (D.18); this results in generally improved agreement between asymptotic predictions for collar speeds and collar speeds in numerical simulations (figure 5a).